# Differentially Private Distributed Inference

### v6

@@ 1 Introduction
In many information environments and information systems, participating agents have privacy and security needs that prevent them from engaging in information exchange at their socially optimal levels (for optimum collective decision making and social learning), e.g., when bound by client confidentiality, obligated by user privacy laws, or need to safeguard their information sources (Acquisti et al. 2015, Bélanger and Crossler 2011, Bélanger and James 2020). In other social and business settings, participants may be reluctant to express unpopular opinions or openly deliberate on controversial or contentious issues. Even in the absence of such privacy, legal, security, or safety concerns, strategic and competitive considerations can lead to suboptimal information sharing and heterogeneous data integration problems between agents (Zhao and Ram 2007). It is therefore reasonable to ask the following research question:
How can agents exchange information to learn from each other despite their privacy needs and security concerns?
@@
The privacy-preserving distributed inference framework in this paper can alleviate the need for complex DUAs and allow centers to communicate privacy-preserving statistics while complying with patient confidentiality and privacy regulations. In 2024, the US Department of Health & Human Services launched the Advancing Clinical Trial Readiness (ACTR) initiative through the ARPA-H Resilient Systems Office (ARPA-H 2024). ACTR emphasizes the development of a decentralized and on-demand clinical trial infrastructure, aiming to make clinical trials accessible to 90% of eligible participants within 30 minutes of their home. A critical aspect of this transformation is automating data extraction and synchronization between electronic health records (EHRs) and case report forms (CRFs), as well as standardizing data collection across diverse centers. Using data from real-world clinical trails, we demonstrate the utility of our privacy-preserving distributed inference framework for multicenter studies.
@@ 2.1 Survival Analysis for Multicenter Clinical Trials
Healthcare centers vary in the demographics of the patients that they serve, and they may even have different target groups (e.g., children’s or women’s hospitals). Although different centers can conduct clinical trials and hypothesis tests on their own, potential data limitations can yield false conclusions, for example, due to small sample sizes or higher prevalence of specific health conditions among some demographics that are under- or overrepresented in their patient population. For this reason, multiple healthcare centers can join multicenter trials to perform hypothesis tests with more accuracy, using larger and more diverse patient samples. To model multicenter trails, we divide patient data from the ACTG study 175 (National Institute of Allergy and Infectious Diseases 1995) equally at random between five centers and use the proportional hazards model (Cox 1972) for survival analysis. For ease of exposition, we consider a simple hypothesis testing scenario to determine the efficacy and safety of a new treatment (ddI) against standard care (ZDV), which are antiretroviral HIV / AIDS medications in the ACTG study 175. For each center $i\in[n]$, we denote their data set of patients by $\bm{S}_{i}$ with cardinality $n_{i}$. The patient $j$ in the center $i$ has the treatment variable $\bm{x}_{ij}\in\{0,1\}$ which indicates if they have received the new treatment (ddI), $\bm{x}_{ij}=1$, or standard care (ZDV), $\bm{x}_{ij}=0$. The effect of the treatment is parametrized by $\theta$. The survival event for each patient is denoted by $\bm{\delta}_{ij}\in\{0,1\}$ with $\bm{\delta}_{ij}=1$ corresponding to death (survival time is observed) and $\bm{\delta}_{ij}=0$ corresponding to survival (survival time is censored). The survival or censoring time is measured in days and is indicated by $\bm{t}_{ij}\in\mathbb{N}$. Following the Cox proportional hazards model, the partial log-likelihood of the patient survival data for center $i$ is given by:
@@
(a) Centralized curves
(b) Curves for one hospital
@@ 3.1 Problem Formulation: Distributed Inference & Learning in Discrete Spaces
A mechanism $\mathcal{M}_{i}:\mathcal{S}\to\mathcal{R}$, acting on private signals, is said to be ${\varepsilon}$-DP with respect to the signal $\bm{s}\in\mathcal{S}$ if and only if for all $X\subseteq\mathcal{R}$ we have that
${\mathbb{P}[\mathcal{M}_{i}(\bm{s})\in X]}\leq e^{\varepsilon}{\mathbb{P}[\mathcal{M}_{i}(\bm{s}^{\prime})\in X]},\text{ for all }\bm{s},\bm{s}^{\prime}\in\mathcal{S},\text{ s.t. }\left\|\bm{s}-\bm{s}^{\prime}\right\|_{1}\leq 1.$
@@
In the multicenter clinical trial example of Section 2.1, the hypothesis space is $\Theta=\{0,\theta_{1}\}$ where $\theta=0$ corresponds to the hypothesis that the two treatments are statistically indistinguishable and $\theta=\theta_{1}<0$ corresponds to the hypothesis that ddI improves patient survival compared to ZDV as is the case based on Figure 3; hence, $\Theta^{\star}=\left\{\theta_{1}\right\}$. The goal of the agents is to collectively infer that $\theta=\theta_{1}$ is the best state (MLE) given the data of the patients because $\Lambda(\theta_{1})>\Lambda(0)$. In Section 4.2, we show how distributed MLE algorithms can be modified for distributed hypothesis testing with statistical significance guarantees.
@@ 3.2 Differentially Private Log-Linear Belief Updates and Opinion Pools
The fastest way to (asymptotically) compute the joint log-likelihood would be for the agents to run a linear consensus algorithm (DeGroot 1974, Olshevsky 2017) in the (privatized) log-likelihood space for each $\theta\in\Theta$. This rule is equivalent to performing log-linear updates (i.e., geometric means) in a (non-Bayesian) belief space (Rahimian and Jadbabaie 2016b). On the other hand, simple DeGroot (linear) averaging of beliefs converges to the average of initial likelihoods (rather than log-likelihoods) and can give biased estimates if used for maximum likelihood estimation. Kayaalp et al. (2023) and (Lalitha et al. 2018, Corollary 2) point out that log-linear learning converges faster than linear learning in the belief space. Marden and Shamma (2012) provide guarantees on the percentage of time that the action profile will be a potential maximizer in potential games using log-linear best-response dynamics.
@@
The ${\varepsilon}$-DP mechanism $\mathcal{M}_{i}$, satisfying Definition 3.1, applied to $\bm{\gamma}_{i}$ produces $\mathcal{M}_{i}(\bm{\gamma}_{i})$ which is privatized belief over the hypothesis space $|\Theta|$. Putting everything together, the expected payoff of agent $i$ at time $t$ is given as (see Eq. (9) in Abbas (2009)):
@@
To implement the mechanism $\mathcal{M}_{i}$, the decision maker draws appropriately chosen zero-mean (w.l.o.g.) noise variables $\bm{d}_{i}(\hat{\theta})$ for each state $\hat{\theta}\in\Theta$ and adds them to the corresponding log-likelihoods $\log\bm{\gamma}_{i}(\hat{\theta})$. Thus, $\mathcal{M}_{i}(\bm{\gamma}_{i})$ can be written as the product of the two distributions $\bm{\gamma}_{i}$ and $\mathfrak{D}_{i}$, where $\mathfrak{D}_{i}(\hat{\theta})=\frac{e^{\bm{d}_{i}(\hat{\theta})}}{\sum_{\bar{\theta}\in\Theta}e^{\bm{d}_{i}(\bar{\theta})}}$. If we write $\bm{\nu}_{i,0}$ as the product of $\bm{\nu}_{i,0}$ together with the uniform measure $\mathcal{U}_{\Theta}$ on $\Theta$ and using the properties of the KL-divergence, we get $D_{KL}\left(\bm{\nu}_{i,0}|\mathcal{M}_{i}(\bm{\gamma}_{i})\right)=D_{KL}\left(\bm{\nu}_{i,0}|\bm{\gamma}_{i}\right)+D_{KL}\left(\mathcal{U}_{\Theta}|\mathfrak{D}_{i}\right)$. The term $D_{KL}\left(\mathcal{U}_{\Theta}|\mathfrak{D}_{i}\right)$ corresponds to the privacy loss and equals
@@
Taking the expectation over $\mathcal{M}_{i}$ we get $\mathbb{E}_{\mathcal{M}_{i}}\left[D_{KL}\left(\mathcal{U}_{\Theta}|\mathfrak{D}_{i}\right)\right]=\mathbb{E}_{\mathcal{M}_{i}}\left[\log\left({\sum_{\bar{\theta}\in\Theta}e^{\bm{d}_{i}(\bar{\theta})}}\right)\right]-\log|\Theta|$. Since the KL divergence is nonnegative, the expected privacy loss with respect to $\mathcal{M}_{i}$ is minimized whenever the noise variables $\bm{d}_{i}(\hat{\theta})$ are independent of each other, and their distribution does not depend on $\hat{\theta}$.
To achieve the aforementioned property, $\bm{d}_{i}(\hat{\theta})$ can be distributed according to the Laplace noise mechanism, and, in addition, the distribution of the noise should be independent of the state $\hat{\theta}\in\Theta$, i.e. $\bm{d}_{i}(\hat{\theta})\sim\mathcal{D}$. The Laplace mechanism possesses further benefits since it minimizes the convergence time of the belief exchange algorithms (cf. Section 4 and prior results by Papachristou and Rahimian (2024) and Koufogiannis et al. (2015)). The global $\ell_{1}$-sensitivity is computed as
$\Delta_{n,\Theta}=\max_{i\in[n],\hat{\theta}\in\Theta}n_{i}g_{i}(\hat{\theta}),\quad\text{where}\quad g_{i}(\hat{\theta})=\max_{\bm{s},\bm{s}^{\prime}\in\mathcal{S}:\|\bm{s}-\bm{s}^{\prime}\|\leq 1}\big{\|}\log\ell_{i}(\bm{s}|\hat{\theta})-\log\ell_{i}(\bm{s}^{\prime}|\hat{\theta})\big{\|}_{1}.$ (6)
In Section 12, we compute (6) for the proportional hazards model. More generally, the signal structure may be such that the sensitivity is unbounded when computed globally over the signal space. In such cases, an additive relaxation of the privacy constraints in Definition 3.1 can be used with a smoothed notion that looks at sensitivity locally at the realized signal values (Nissim et al. 2007).
The best update rule that optimizes the time-varying utility of Equation 5 yields the log-linear update rules (Abbas 2009, Proposition 4). The above can be extended to the online setting where data $\bm{S}_{i,t}$ with $n_{i,t}\geq 0$ data points for each agent $i$ are intermittently arriving at each iteration $t$, by introducing a time-varying utility and measuring the likelihood at each iteration $t$. In that regime, first, the agents form the probability measure $\bm{\gamma}_{i,t}$ such that $\bm{\gamma}_{i,t}(\hat{\theta})=\frac{\ell_{i}(\bm{S}_{i,t}|\hat{\theta})}{\sum_{\bar{\theta}\in\Theta}\ell_{i}(\bm{S}_{i,t}|\bar{\theta})}$ for all $\hat{\theta}\in\Theta$ (when $n_{i,t}=0$ we set $\ell_{i}(\bm{S}_{i,t}|\hat{\theta})=1$ for all $\hat{\theta}\in\Theta$). Then the agents have the time-varying utility:
@@ 4.1 Differentially Private Distributed MLE
The distributed MLE algorithm is repeated $K$ times for each of the $|\Theta|$ states, and an attacker would have more observations of the same signals; therefore, we need to rescale the privacy guarantee for each run and each state by $K|\Theta|$. Thus, choosing $\mathcal{D}_{i}({\varepsilon})=\mathrm{Lap}\left(\frac{K|\Theta|\Delta_{n,\Theta}}{{\varepsilon}}\right)$ satisfies ${\varepsilon}$-DP and the assumptions of Theorem 4.1. Our nonasymptotic analysis (presented next) shows that $\mathcal{D}_{i}({\varepsilon})=\mathrm{Lap}\left(\frac{K|\Theta|\Delta_{n,\Theta}}{{\varepsilon}}\right)$ is indeed optimal to minimize the required number of iterations per round.
@@
Recall that $\Gamma_{n,\Theta}$ and $l_{n,\Theta}$ that appear in the communication complexity bounds in Theorem 4.2 are properties of the signal structure defined in Equations 8 and 9). To minimize communication complexity, it suffices to pick the optimal threshold that makes the terms inside the maximum equal – and, hence, the maximum is minimized – which corresponds to $\varrho^{\mathrm{AM}}_{\star}=\left(\frac{|\Theta|^{2}(n-1)K(n\Gamma_{n,\Theta}+V_{n,\Theta})}{2\log(1/(1-\beta))}\right)^{\frac{\log 2}{\log(2/a_{n}^{\star})}}\left(\frac{l_{n,\Theta}}{2n}\right)^{\frac{-\log a_{n}^{\star}}{\log(2/a_{n}^{\star})}},$ for AM and $\varrho^{\mathrm{GM}}_{\star}=\left(\frac{|\Theta|^{2}(n-1)(n\Gamma_{n,\Theta}+V_{n,\Theta})}{2\alpha\sqrt{K}}\right)^{\frac{\log 2}{\log(2/a_{n}^{\star})}}\left(\frac{l_{n,\Theta}}{2n}\right)^{\frac{-\log a_{n}^{\star}}{\log(2/a_{n}^{\star})}}$ for GM.
@@
If the density of MLE states $f^{\star}$ satisfies
@@
the thresholds are set equal, i.e., $\tau^{\mathrm{thres},2}=\tau^{\mathrm{thres},1}$ and ${\hat{\tau}}^{\mathrm{thres},1}={\hat{\tau}}^{\mathrm{thres},2}$ which corresponds to $\hat{\Theta}^{\mathrm{thres},2}_{i,T}=\hat{\Theta}^{\mathrm{thres},1}_{i,T}$ and $\pi_{1}=(1-\pi_{2})\frac{p_{2}}{p_{1}}-1$ for some $\pi_{2}>0$, and $K\geq\max\left\{\frac{\log(|\Theta|/\alpha)}{2\pi_{1}^{2}},\frac{\log(|\Theta|/(1-\beta))}{2\pi_{2}^{2}}\right\}$ we have that $\lim_{T\to\infty}\mathbb{P}\left[\hat{\Theta}^{\mathrm{thres},2}_{i,T}=\Theta^{\star}\right]\geq\beta-\alpha.$

### v7

@@ 1 Introduction
In many information environments and information systems, participating agents have privacy and security needs that prevent them from engaging in information exchange at their socially optimal levels (for optimum collective decision making and social learning), e.g., when bound by client confidentiality, obligated by user privacy laws, or when they need to safeguard their information sources (Acquisti et al. 2015, Bélanger and Crossler 2011, Bélanger and James 2020). In other social and business settings, participants may be reluctant to express unpopular opinions or openly deliberate on controversial or contentious issues. Even in the absence of such privacy, legal, security, or safety concerns, strategic and competitive considerations can lead to suboptimal information sharing and heterogeneous data integration problems between agents (Zhao and Ram 2007). It is therefore reasonable to ask the following research question:
How can agents exchange information for collective inference and learn from each other despite their privacy needs and security concerns?
@@
The privacy-preserving distributed inference framework in this paper can alleviate the need for complex DUAs and allow centers to communicate privacy-preserving statistics while complying with patient confidentiality and privacy regulations. In 2024, the US Department of Health & Human Services launched the Advancing Clinical Trial Readiness (ACTR) initiative through the ARPA-H Resilient Systems Office (ARPA-H 2024). ACTR emphasizes the development of a decentralized and on-demand clinical trial infrastructure, aiming to make clinical trials accessible to 90% of eligible participants within 30 minutes of their home. A critical aspect of this transformation is automating data extraction and synchronization between electronic health records (EHRs) and case report forms (CRFs), as well as standardizing data collection across diverse centers. Using data from real-world clinical trials, we demonstrate the utility of our privacy-preserving distributed inference framework for multicenter studies.
@@ 2.1 Survival Analysis for Multicenter Clinical Trials
Healthcare centers vary in the demographics of the patients that they serve, and they may even have different target groups (e.g., children’s or women’s hospitals). Although different centers can conduct clinical trials and hypothesis tests on their own, potential data limitations can yield false conclusions, for example, due to small sample sizes or higher prevalence of specific health conditions among some demographics that are under- or overrepresented in their patient population. For this reason, multiple healthcare centers can join multicenter trials to perform hypothesis tests with more accuracy, using larger and more diverse patient samples. To model multicenter trials, we divide patient data from the ACTG study 175 (National Institute of Allergy and Infectious Diseases 1995) equally at random between five centers and use the proportional hazards model (Cox 1972) for survival analysis. For ease of exposition, we consider a simple hypothesis testing scenario to determine the efficacy and safety of a new treatment (ddI) against standard care (ZDV), which are antiretroviral HIV / AIDS medications in the ACTG study 175. For each center $i\in[n]$, we denote their data set of patients by $\bm{S}_{i}$ with cardinality $n_{i}$. The patient $j$ in the center $i$ has the treatment variable $\bm{x}_{ij}\in\{0,1\}$ which indicates if they have received the new treatment (ddI), $\bm{x}_{ij}=1$, or standard care (ZDV), $\bm{x}_{ij}=0$. The effect of the treatment is parametrized by $\theta$. The survival event for each patient is denoted by $\bm{\delta}_{ij}\in\{0,1\}$ with $\bm{\delta}_{ij}=1$ corresponding to death (survival time is observed) and $\bm{\delta}_{ij}=0$ corresponding to survival (survival time is censored). The survival or censoring time is measured in days and is indicated by $\bm{t}_{ij}\in\mathbb{N}$. Following the Cox proportional hazards model, the partial log-likelihood of the patient survival data for center $i$ is given by:
@@
(a) Centralized
(b) One hospital
@@ 3.1 Problem Formulation: Distributed Inference & Learning in Discrete Spaces
A mechanism $\mathcal{M}_{i}:\mathcal{S}\to\mathcal{R}$, acting on private signals, is said to be ${\varepsilon}$-DP with respect to the signal $\bm{s}\in\mathcal{S}\subset\mathbb{R}^{d}$ if and only if for all $X\subseteq\mathcal{R}$ we have that
${\mathbb{P}[\mathcal{M}_{i}(\bm{s})\in X]}\leq e^{\varepsilon}{\mathbb{P}[\mathcal{M}_{i}(\bm{s}^{\prime})\in X]},\text{ for all }\bm{s},\bm{s}^{\prime}\in\mathcal{S},\text{ s.t. }\left\|\bm{s}-\bm{s}^{\prime}\right\|_{1}\leq 1.$ (4)
@@
In the multicenter clinical trial example of Section 2.1, the state space is $\Theta=\{0,\theta_{1}\}$ where $\theta=0$ corresponds to the hypothesis that the two treatments are statistically indistinguishable and $\theta=\theta_{1}<0$ corresponds to the hypothesis that ddI improves patient survival compared to ZDV as is the case based on Figure 3; hence, $\Theta^{\star}=\left\{\theta_{1}\right\}$. The goal of the agents is to collectively infer that $\theta=\theta_{1}$ is the best state (MLE) given the data of the patients because $\Lambda(\theta_{1})>\Lambda(0)$. In Section 4.2, we show how distributed MLE algorithms can be modified for distributed hypothesis testing with statistical significance guarantees.
@@ 3.2 Differentially Private Log-Linear Belief Updates and Opinion Pools
The fastest way to (asymptotically) compute the joint log-likelihood would be for the agents to run a linear consensus algorithm (DeGroot 1974, Olshevsky 2017) in the (privatized) log-likelihood space for each $\theta\in\Theta$. This rule is equivalent to performing log-linear updates (i.e., geometric means) in a (non-Bayesian) belief space (Rahimian and Jadbabaie 2016b). On the other hand, simple DeGroot (linear) averaging of beliefs converges to the average of initial likelihoods (rather than log-likelihoods) and can give biased estimates if used for maximum likelihood estimation. Kayaalp et al. (2023) and Lalitha et al. (2018, Corollary 2) point out that log-linear learning converges faster than linear learning in the belief space. Marden and Shamma (2012) provide guarantees on the percentage of time that the action profile will be a potential maximizer in potential games using log-linear best-response dynamics.
@@
The ${\varepsilon}$-DP mechanism $\mathcal{M}_{i}$, satisfying Definition 3.1, applied to $\bm{\gamma}_{i}$ produces $\mathcal{M}_{i}(\bm{\gamma}_{i})$ which is the privatized belief over the state space $|\Theta|$. Putting everything together, the expected payoff of agent $i$ at time $t$ is given as (see Eq. (9) in Abbas (2009)):
@@
To implement the mechanism $\mathcal{M}_{i}$, the decision maker draws appropriately chosen, without loss of generality, zero-mean noise variables $\bm{d}_{i}(\hat{\theta})$ for each state $\hat{\theta}\in\Theta$ and adds them to the corresponding log-likelihoods $\log\bm{\gamma}_{i}(\hat{\theta})$. Thus, $\mathcal{M}_{i}(\bm{\gamma}_{i})$ can be written as the product of the two distributions $\bm{\gamma}_{i}$ and $\mathfrak{D}_{i}$, where $\mathfrak{D}_{i}(\hat{\theta})=\frac{e^{\bm{d}_{i}(\hat{\theta})}}{\sum_{\bar{\theta}\in\Theta}e^{\bm{d}_{i}(\bar{\theta})}}$. If we write $\bm{\nu}_{i,0}$ as the product of $\bm{\nu}_{i,0}$ and the uniform distribution $\mathcal{U}_{\Theta}$ on $\Theta$, then we can use the properties of the KL divergence to write $D_{KL}\left(\bm{\nu}_{i,0}|\mathcal{M}_{i}(\bm{\gamma}_{i})\right)=D_{KL}\left(\bm{\nu}_{i,0}|\bm{\gamma}_{i}\right)+D_{KL}\left(\mathcal{U}_{\Theta}|\mathfrak{D}_{i}\right)$. The term $D_{KL}\left(\mathcal{U}_{\Theta}|\mathfrak{D}_{i}\right)$ corresponds to the privacy loss for the utility in Equation 6 and equals
@@
Taking the expectation over $\mathcal{M}_{i}$ we get $\mathbb{E}_{\mathcal{M}_{i}}\left[D_{KL}\left(\mathcal{U}_{\Theta}|\mathfrak{D}_{i}\right)\right]=\mathbb{E}_{\mathcal{M}_{i}}\left[\log\left({\sum_{\bar{\theta}\in\Theta}e^{\bm{d}_{i}(\bar{\theta})}}\right)\right]-\log|\Theta|$. Since the KL divergence is nonnegative, the expected privacy loss with respect to $\mathcal{M}_{i}$ is minimized whenever the noise variables $\bm{d}_{i}(\hat{\theta})$ are independent of each other and their distribution does not depend on $\hat{\theta}$, i.e., $\bm{d}_{i}(\hat{\theta})=\bm{d}_{i}$.
To satisfy the ${\varepsilon}$-DP requirement in Definition 3.1, $\bm{d}_{i}$ can be set according to the Laplace mechanism (Dwork and Roth 2014, Section 3.3), which adds Laplace distribution noise in the $\log$ space: $\bm{d}_{i}\sim\mathcal{D}=\mathrm{Lap}\left(\frac{|\Theta|\Delta_{n,\Theta}}{{\varepsilon}}\right)$, where ${|\Theta|\Delta_{n,\Theta}}/{{\varepsilon}}$ is the scale parameter of the Laplace distribution and the global $\ell_{1}$-sensitivity, $\Delta_{n,\Theta}$, is defined as:
$\Delta_{n,\Theta}=\max_{i\in[n],\hat{\theta}\in\Theta}n_{i}g_{i}(\hat{\theta}),\quad\text{where}\quad g_{i}(\hat{\theta})=\max_{\bm{S}_{i},\bm{S}^{\prime}_{i}\in\mathcal{S}:\|\bm{S}_{i}-\bm{S}^{\prime}_{i}\|\leq 1}\big{\|}\log\ell_{i}(\bm{S}_{i}|\hat{\theta})-\log\ell_{i}(\bm{S}^{\prime}_{i}|\hat{\theta})\big{\|}_{1}.$ (7)
Intuitively, when $\log$-likelihoods are highly sensitive to signal values more noise is required to mask the input signals at the ${\varepsilon}$-DP level. Here we have used ${\varepsilon}/|\Theta|$ privacy budget per state to ensure that $\bm{\nu}_{i,0}\in\mathrm{Simplex}(\Theta)$ is ${\varepsilon}$-DP overall by the composition property of DP (Dwork and Roth 2014, Theorem 3.14), after computing its value at all states $\hat{\theta}\in\Theta$. In Section 12, we bound (7) for the proportional hazards model. More generally, the signal structure may be such that the sensitivity is unbounded when computed globally over the signal space. In such cases, an additive relaxation of the privacy constraints in Definition 3.1 can be used with a smoothed notion that looks at sensitivity locally at the realized signal values (Nissim et al. 2007). The Laplace mechanism possesses further benefits since it minimizes the convergence time of the belief exchange algorithms (cf. Section 4 and prior results by Papachristou and Rahimian (2024) and Koufogiannis et al. (2015)). The best update rule that optimizes the time-varying utility of Equation 6 yields the log-linear update rules (Abbas 2009, Proposition 4). The above can be extended to the online setting where data $\bm{S}_{i,t}$ with $n_{i,t}\geq 0$ data points for each agent $i$ are intermittently arriving at each iteration $t$, by introducing a time-varying utility and measuring the likelihood at each iteration $t$. In that regime, first, the agents form the probability measure $\bm{\gamma}_{i,t}$ such that $\bm{\gamma}_{i,t}(\hat{\theta})=\frac{\ell_{i}(\bm{S}_{i,t}|\hat{\theta})}{\sum_{\bar{\theta}\in\Theta}\ell_{i}(\bm{S}_{i,t}|\bar{\theta})}$ for all $\hat{\theta}\in\Theta$ (when $n_{i,t}=0$ we set $\ell_{i}(\bm{S}_{i,t}|\hat{\theta})=1$ for all $\hat{\theta}\in\Theta$). Then the agents have the time-varying utility:
@@ 4.1 Differentially Private Distributed MLE
The distributed MLE algorithm is repeated $K$ times for each of the $|\Theta|$ states, and an attacker would have more observations of the outputs from the same signals; therefore, we need to rescale the privacy budget by $K|\Theta|$ to account for the number of runs and states. Thus, choosing $\mathcal{D}_{i}({\varepsilon})=\mathrm{Lap}\left(\frac{K|\Theta|\Delta_{n,\Theta}}{{\varepsilon}}\right)$ satisfies ${\varepsilon}$-DP and the assumptions of Theorem 4.1. Our nonasymptotic analysis (presented next) shows that $\mathcal{D}_{i}({\varepsilon})=\mathrm{Lap}\left(\frac{K|\Theta|\Delta_{n,\Theta}}{{\varepsilon}}\right)$ is indeed optimal to minimize the required number of iterations per round.
@@
Recall that $\Gamma_{n,\Theta}$ and $l_{n,\Theta}$ that appear in the communication complexity bounds in Theorem 4.2 are properties of the signal structure defined in Equations 9 and 10). To minimize communication complexity, it suffices to pick the optimal threshold that makes the terms inside the maximum equal – and, hence, the maximum is minimized – which corresponds to $\varrho^{\mathrm{AM}}_{\star}=\left(\frac{|\Theta|^{2}(n-1)K(n\Gamma_{n,\Theta}+V_{n,\Theta})}{2\log(1/(1-\beta))}\right)^{\frac{\log 2}{\log(2/a_{n}^{\star})}}\left(\frac{l_{n,\Theta}}{2n}\right)^{\frac{-\log a_{n}^{\star}}{\log(2/a_{n}^{\star})}},$ for AM and $\varrho^{\mathrm{GM}}_{\star}=\left(\frac{|\Theta|^{2}(n-1)(n\Gamma_{n,\Theta}+V_{n,\Theta})}{2\alpha\sqrt{K}}\right)^{\frac{\log 2}{\log(2/a_{n}^{\star})}}\left(\frac{l_{n,\Theta}}{2n}\right)^{\frac{-\log a_{n}^{\star}}{\log(2/a_{n}^{\star})}}$ for GM.
@@
If the density of MLE states $f^{\star}$ satisfies:
@@
the thresholds are set equal, i.e., $\tau^{\mathrm{thres},2}=\tau^{\mathrm{thres},1}$ and ${\hat{\tau}}^{\mathrm{thres},1}={\hat{\tau}}^{\mathrm{thres},2}$ which corresponds to $\hat{\Theta}^{\mathrm{thres},2}_{i,T}=\hat{\Theta}^{\mathrm{thres},1}_{i,T}$ and $\pi_{1}=(1-\pi_{2})\frac{p_{2}}{p_{1}}-1$ for some $\pi_{2}>0$, and $K\geq\max\left\{\frac{\log(|\Theta|/\alpha)}{2\pi_{1}^{2}},\frac{\log(|\Theta|/(1-\beta))}{2\pi_{2}^{2}}\right\}$, then we have that $\lim_{T\to\infty}\mathbb{P}\left[\hat{\Theta}^{\mathrm{thres},2}_{i,T}=\Theta^{\star}\right]\geq\beta-\alpha.$